hi, i am a mature student am have got the inequalities so far but am stuck on how to solve squared ones i.e. x squared etc. If you could help that'd be great.
regards denise

Answer by checkley71(8403)

THIS MIGHT HELP X~2 IS THE SAME AS X*X OR X TIMES X
3~2 IS THE SAME AS 3*3 OR 9
X~2+4X+4=0 THEN (X+2)(X+2)=0 OR X+2=0 OR X=-2 AND X+2=0 X=-2